Is there a difference between potential and potential energy? Why?
I don't understand completely. I think there is a difference and that they have different units. Could you explain?  
Answer #1

Electric potential is the potential energy per unit charge, just as electric field is the force per unit charge. It's a way to describe the system independent of the "test charge".

So if it takes 10 eV of energy to move a single elementary charge (like a proton) from A to B, we say the difference in the proton's potential energy is 10 eV between those points. But independent of the proton, we say the difference in potential between those two points is 10 Volts.

The charge we're moving has potential energy U. The field we're moving it through has potential. V = U/q, or U = qV.
Answer #2
potential is the energy of a unit charge and potential energy is the energy of the charge... potential is denoted by V and energy is denoted by U ==> U=q*V
Answer #3

Potential Energy

W =
Answer #4

Electric potential at a point in an Electric field (for instance, let's take it as uniform) is the amount of work done in moving a unit positive charge from infinity to that point against electric forces. (attraction or repulsion)

whereas Electric potential difference between two points in an Electric field is defined as the amount of work done in moving unit positive charge from one point to another point against the electric force.

In other words, Potential difference is simply the difference of scalar potentials between two points or anything you consider.
Electric potential at a point is given by

V=q/4??r

But, Potential energy requires the magnitude of another charge also... If r is the distance of separation of two charges, P.E. is given by

Ep=qQ/4??r.

In other words, Ep= VQ